What sum are you least likely to get rolling two dice?
all numbers on a die have the same probability to land so any number between 2 and 12 are just as likely as any other number between 2 and 12, just as any other die hase the same probability 2 d20, 2-40 any to are just as likely as the rest. so 0 is the least likely number.
Well, yes, since 0 is impossible to be the sum of 2 dice, it is least likely. So is any other number less than 2 or greater than 12, or a number that is not a whole number. But of those that actually can come up as the sum of 2 dice, 7 is the most likely because there are 6 ways to get it (1 and 6, 2 and 5, 3 and 4, 4 and 3, etc.). The least likely sums are 2 and 12, because each can only occur 1 way (a pair of 1's or a pair of 6's). These are referred 2 by gamblers as snake eyes and box cars.
What is the probability of getting 2 heads when you toss three coins?
If you mean 'at least' 2 heads, the probability is 50%. If you mean exactly 2,
the probability is 3/8, or 37.5%.
There are 3 independent coin tosses, each of which is equally likely to come up
heads or tails. That's a total of 2 * 2 * 2 or 8 possible outcomes (HHH, HHT, HTH,
etc.). Of these, 4 include 2 or 3 heads, which is half of 8. Only 3 include exactly 2
heads, so the probability of that is 3/8.
What is the probability of drawing a ten and a spade?
If you draw enough cards then it is very close to 1 - almost a certainty.
What is the probability of being dealt 4 clubs from a deck of 52 cards?
13/52*12/51*11/50*10/49
= 1/4*4/7*11/50*10/49
= 440/68600
= 44/6860
= 11/1715
Which one is more convienent mutually exclusive or non mutually?
It must be "mutually exclusive" since "non mutually" does not even mean anything!
What is the theoretical probability of not rolling a 2?
1/6= 2 because there is only one 2.
Therefore the theoretical probability of not rolling a two is the same as everything but two so 5/6.
How many 5 are in a deck if cards?
4 cards has 5 in it. 2 "5" on each card. means totally 8 of "5" on cards
Risk
How many permutations are in the word peak?
The first letter of the perm can be any of the 4, the second any of the remaining 3, the third either of the remaining 2. There are thus 4 x 3 x 2 ie 24 permutations.
What number can be used to describe a certain event?
The number one can be used to describe a certain event.
What is the probability of 2 coins landing on heads?
well since the coins have two sides,there is a 50% chance of it landing on heads
How many permutations are in the word geometry?
geometry has 8 letters, 2 of which are the same (e)
So, the answer is 8!/2! = 20,160
I am assuming that this die is fair die and the coin is not biased. The probability of getting a number less than 3 is the probability of rolling a 1 or a 2 i.e. 1/6 + 1/6 = 2/6 which simplifies to 1/3. The probability of getting a head when you flip a fair coin is 1/2. Both are independent events, so the probability of getting a number less than 3 and getting a head is 1/3 x 1/2 = 1/6. One can get the same answer from a sample space diagram
Is the distribution of sample means always a normal distribution If not why?
It need not be if:
the number of samples is small;
the elements within each sample, and the samples themselves are not selected independently.
What is the probability that you toss a head or a tail?
Assuming a two-sided coin, and that you make the the toss, the probability of tossing a head or a tail is 100%. The probability of tossing a head is 50%. The probability of tossing a tail is 50%.
What is the probability of drawing a red king from a deck of 52?
Probability of getting a King is 4/52 = 1/13. Having drawn the King, for it to be a Red King, the probability is 2/4 = 1/2. Hence the conditional probability is (1/13)x(1/2) = 1/26
What is the probability of spinning an even number all four times?
That depends on the numbers you can spin.